### v1

# Hypercube-RAG: Hypercube-Based Retrieval-Augmented Generation for In-domain Scientific Question-Answering

Abstract
Large language models (LLMs) often need to incorporate external knowledge to solve theme-specific problems. Retrieval-augmented generation (RAG), which empowers LLMs to generate more qualified responses with retrieved external data and knowledge, has shown its high promise. However, traditional semantic similarity-based RAGs struggle to return concise yet highly relevant information for domain knowledge-intensive tasks, such as scientific question-answering (QA). Built on a multi-dimensional (cube) structure called Hypercube, which can index documents in an application-driven, human-defined, multi-dimensional space, we introduce the Hypercube-RAG, a novel RAG framework for precise and efficient retrieval. Given a query, Hypercube-RAG first decomposes it based on its entities and topics and then retrieves relevant documents from cubes by aligning these decomposed components with hypercube dimensions. Experiments on three in-domain scientific QA datasets demonstrate that our method improves accuracy by and boosts retrieval efficiency by , measured as relative gains over the strongest RAG baseline. More importantly, our Hypercube-RAG inherently offers explainability by revealing the underlying predefined hypercube dimensions used for retrieval. The code and data sets are available at https://github.com/JimengShi/Hypercube-RAG.
Hypercube-RAG: Hypercube-Based Retrieval-Augmented Generation
for In-domain Scientific Question-Answering
Jimeng Shi1, Sizhe Zhou2, Bowen Jin2, Wei Hu2, Shaowen Wang2, Giri Narasimhan1, Jiawei Han2 1Florida International University 2University of Illinois Urbana-Champaign {jshi008,giri}@fiu.edu, {sizhez,bowenj4,weih9,shaowen,hanj}@illinois.edu
1 Introduction
Large language models (LLMs) often suffer from hallucinations and factual inaccuracies, especially for in-domain scientific question-answering. To address this phenomenon, retrieval-augmented generation (RAG) has emerged as the de facto approach Huang et al. (2025), incorporating external domain knowledge to generate contextually relevant responses Huang et al. (2025). Despite recent advancements, they exhibit key limitations in theme-specific applications on which the fine-grained topics are highly focused. For instance, retrieved documents are often semantically similar to the query but overlook specialized themes (e.g., terminology or nuanced contextual cues) prevalent in specific literature or reports Ding et al. (2024). Moreover, the retrieval process often suffers from inefficiency and limited transparency, which are critical in high-stakes domains, such as environment Wang et al. (2025), traffic Hussien et al. (2025), and healthcare Pujari et al. (2023).
The conventional sparse lexical retrieval ranks a set of documents based on the appearance of query terms Kadhim (2019). Due to its reliance on the exact token overlap, it precisely captures specific themes or topics, and the retrieval process is efficient and interpretable. However, it struggles to retrieve contextually related or paraphrased documents that require semantic understanding. To mitigate the above limitation, dense embedding retrievers were proposed by computing the similarity between vector embeddings of query-document pairs Karpukhin et al. (2020). While contextual understanding is relatively improved, the hurdle of missing specific themes remains, causing semantically similar but off-topic retrievals Krishna et al. (2021); Kang et al. (2024). Figure 1 shows an example where the Semantic-RAG returns documents related to precipitation in Florida, but overlooks the critical location-specific information for ‘‘Melbourne Beach’’. Moreover, retrieval in an embedding space makes it challenging to interpret why certain documents were selected Ji et al. (2019). Additionally, a graph-based RAG represents documents as graph structures, where nodes correspond to entities or concepts and edges capture their relationships Peng et al. (2024). They offer explanations by outputting the traversed subgraph. However, traversing such subgraphs introduces a significant risk of information overload, as the expanding neighborhoods may include substantial amounts of irrelevant themes Wang et al. (2024b). Graph traversal also poses significant scalability bottlenecks due to its inefficient computational complexity of , where and are the number of nodes and edges in the traversed subgraph Devezas and Nunes (2021).
The methods discussed above tend to perform poorly in at least one of the aspects among accuracy, efficiency, and explainability. This motivates us to develop a RAG system that overcomes this weakness. In this pursuit, we identified the text cube as a promising technique Tao et al. (2018); Wang et al. (2023). A text cube is an inherently explainable multidimensional structure that allocates documents into cubes along various human-defined dimensions, such as location, date, event, and theme. Each dimension is populated with fine-grained labels (e.g., ‘Miami Downtown’, ‘2021’, ‘Hurricane Irma’, ‘flooding’) extracted from documents. Relevant documents can be easily indexed by these labels represented in cubes. Furthermore, the explicit theme dimension makes the text cube particularly well-suited for theme-specific applications where users Tao et al. (2018) primarily focus on fine-grained inquiries, e.g., where, when, and which hurricane events hit or how much rainfall hurricanes result in certain areas.
In this work, we implement a multi-dimensional (cube) structure called Hypercube, which indexes documents to cubes based on fine-grained labels in a human-defined, multi-dimensional space. Building on this structure, we propose Hypercube-RAG, a novel RAG framework to enhance the retrieval process by retrieving relevant documents from the corresponding cube cells. Hypercube-RAG exhibits three characteristics simultaneously: (1) Efficiency. The retrieval process quickly narrows down the search to the right cube cell(s) based on the labels with the constant time , while taking to fetch documents assigned in that cell. (2) Explainability. Hypercube is inherently explainable as fine-grained labels in cubes represent the compact information in documents (see Figure 2, right and Figure 3); (3) Accuracy. The cube label-based search supports both sparse and dense embedding strategies, effective for capturing uncommon thematic terminology and semantically relevant information (Section 3.3). Overall, our contributions are summarized as follows:
We identify the shortcomings of conventional RAG methods, especially for theme-specific question-answering in scientific domains.
We propose Hypercube-RAG, a simple yet accurate, efficient, and explainable RAG method.
We conduct experiments on three datasets, demonstrating that our Hypercube-RAG outperforms other baseline methods in accuracy and efficiency when it comes to in-domain question-answering.
2 Related Work
2.1 Text-based RAG
BM25 Robertson and Walker (1994); Trotman et al. (2014) is a classical sparse retrieval method based on TF-IDF principles, widely used in information retrieval. It ranks documents by scoring query term matches, with adjustments for term frequency saturation and document length. While efficient and interpretable, BM25 lacks semantic understanding and often misses paraphrased or contextually related documents without exact lexical overlap, limiting its effectiveness in tasks requiring deeper language comprehension. Dense embedding retrievers that compute contextual similarity between the given query and a document/chunk in the embedding space Jiang et al. (2023). Common retrievers consist of DPR Karpukhin et al. (2020), Contriever Izacard et al. (2022), e5 Wang et al. (2022), and ANCE Xiong et al. (2021). By capturing deeper semantic relationships, these methods improve performance in open-domain question answering and generative tasks, especially when surface-form matching fails. However, dense retrieval can return semantically close yet topically irrelevant documents, introducing noise and hallucinations. Moreover, its opaque nature also hinders interpretability Ji et al. (2019), limiting adoption in high-stakes or scientific domains. Recent efforts aim to mitigate these limitations by incorporating hybrid retrieval techniques Maillard et al. (2021) and domain-adaptive retrievers Lee et al. (2021).
2.2 Structured RAG
Graph-based RAG methods enhance the standard RAG paradigm by introducing structured knowledge representations, such as entity or document graphs. GraphRAG Edge et al. (2024) derives an entity knowledge graph from the source documents and gathers summaries for all groups of closely related entities. LightRAG Guo et al. (2024) employs a dual-level retrieval system that enhances comprehensive information retrieval from both low-level and high-level knowledge discovery. HippoRAG Gutiérrez et al. (2024) is a knowledge graph-based retrieval framework inspired by the hippocampal indexing theory of human long-term memory. HippoRAG 2 Gutiérrez et al. (2025) builds on top of HippoRAG and enhances it with deeper passage integration and more effective online use of an LLM. SELF-RAG Asai et al. (2023) enhances an LLM’s quality and factuality through retrieval and self-reflection. RA-DIT Lin et al. (2023) fine-tunes the pre-trained LLMs or retrievers to incorporate more up-to-date and relevant knowledge. Despite these improvements, graph construction and reasoning can be computationally intensive due to massive nodes and edges.
Text Cube.
Wang et al. (2023) designs the text cube to structure spatial-related data across geospatial dimensions, thematic categories, and diverse application semantics. STREAMCUBE implements the text cube to incorporate both spatial and temporal hierarchies Feng et al. (2015). Doc2Cube Tao et al. (2018) automates the allocation of documents into a text cube to support multidimensional text analytics. Despite these elaborate designs, the text cube has not yet been explored in the context of retrieval-augmented generation (RAG) systems.
3 Hypercube-RAG
Figure 4 illustrates the Hypercube-RAG framework through a scientific question-answering example. In the following, we define the hypercube structure (Section 3.1), describe the process for constructing the hypercube (Section 3.2), explain how two sparse and dense embedding retrieval strategies are combined inside of hypercube (Section 3.3), and present how to rank documents retrieved using hypercube (Section 3.4).
3.1 Hypercube Formulation
Given a text corpus with documents: , the hypercube is designed as a multidimensional data structure, , where is the number of hypercube dimensions, and is the dimension. For any document , a text classification algorithm (e.g, TeleClass Zhang et al. (2025)) will assign it to one or more cube cells. This process is equivalent to assigning a -dimensional labels () for a document, , where label represents the category of in dimension .
3.2 Hypercube Construction
For most scientific applications, theme, location, date, event, and/or person are often of high interest. The core idea of the hypercube is to identify fine-grained entities or topics contained in the corpus and index documents in cubes with the appropriate dimensions, thereby facilitating the subsequent retrieval. In this work, we employ multiple methods to construct a holistic and precise knowledge hypercube.
Name Entity Recognition (NER).
We first pre-define hypercube dimensions LOCATION, DATE, EVENT, PERSON and ORGANIZATION. For each dimension, we utilize the pretrained language models (PLMs) Devlin et al. (2019); Liu et al. (2019) to extract salient entities from documents: , where is the set of named entities identified in the document, , and refers to a PLM. For instance, those entities belonging to the EVENT dimension are represented:
Theme-specific Extraction.
While the above entities can be captured by NER models, abstract and intricate theme-specific information may be missed, such as ‘hurricane path’, ‘rainfall intensity’, and ‘climate change’. Therefore, we add another hypercube dimension THEME, such key theme-specific phrases can be extracted. We implement it with KeyBERT111https://github.com/MaartenGr/KeyBERT?tab=readme-ov-file, a language model pre-trained on large-scale text corpora, which demonstrates strong performance in capturing general topical relevance.
Using the methods described above, we extract fine-grained entities, topics, and phrases, and assign them to different hypercube dimensions, including LOCATION, DATE, EVENT, ORGANIZATION, PERSON, and THEME. To enhance coverage of in-domain semantics, we incorporate human-in-the-loop curation to expand the THEME dimension, ensuring it captures a comprehensive set of relevant topics. Additional implementation details are provided in Appendix E.
3.3 Hypercube Retrieval
A constructed knowledge hypercube can serve for the retrieval process based on its fine-grained labels. Given a query , an LLM first decomposes it into a set of entities and topics, , aligned with various dimensions (see the prompt in Figure 4), then retrieval inside of hypercube is performed by matching these decomposed components with cube labels. Our hypercube supports two matching strategies: sparse exact lexical match and dense embedding match. We prioritize leveraging the exact matching strategy due to its high precision, scoring documents based on the frequency with which query entities appear within them:
where is an indicator function that equals 1 if an entity of the query exactly matches a term in the document, and 0 otherwise.
Considering the example in Figure 4, the last two components from the query, Melbourne Beach and Tropical Storm Fay, are handled using the exact matching strategy, as they correspond precisely to fine-grained labels in the hypercube. However, the first component, Rainfall, does not exactly match any label(s) in the hypercube, even though its matched label should be Rain.
To address such cases, Hypercube-RAG supports semantic retrieval by computing the similarity between query components and fine-grained cube labels. When the similarity score exceeds a predefined threshold , semantic retrieval is triggered. Specifically, both the query entities and cube labels are projected into an embedding space, , with an encoder. It can be represented as:
where refers to the similarity function.
3.4 Hypercube Ranking
The retrieved documents need to be provided to the LLM as contextual input. However, long-context RAG methods do not always improve the quality of LLM responses, since they may introduce irrelevant or noisy information Liu et al. (2024); Jin et al. (2025). This raises a critical challenge: how to precisely filter and select theme-relevant documents. To address this, we prioritize returning those documents that fully cover key components derived from the query:
If no documents fully cover all query components (i.e., ), we return the next best set of documents with the highest partial coverage:
4 Experiments
4.1 Datasets
In this work, we focus on studying Hypercube-RAG within the context of a single hypercube configuration. While more general domains may benefit from modeling multiple hypercubes, our evaluation is centered on one scientific domain to enable a focused and controlled analysis. We also provide a more detailed discussion in Appendix C. We collected the text corpus from a subset of the Science Daily Climate Change dataset (SciDCC) Mishra and Mittal (2021), which was created by web scraping from the “Geography” and “Climate” topics in the environmental science section of the Science Daily website222https://www.sciencedaily.com/. We created an aging dam corpus by collecting news articles related to dam failures in the United States from Google News333https://news.google.com/. Then, an LLM (GPT-4o) is employed to generate 300, 300, and 186 QA pairs, respectively. Each pair was manually validated to ensure that the corresponding question could be answered using information contained in the corpus. To ensure representativeness, each dataset generates three types of QA pairs: (a) long-form, free-style; (b) short-term factual; and (c) quantitatively factual. We present the summary in Table 1 and more details in Appendix B.
| Dataset | # Doc. | Doc. Length | # QA | Ques. Length | Ans. Length |
| Hurricane | 844 | 521781 | 300 | 627 | 2119 |
| Geography | 432 | 221646 | 300 | 534 | 279 |
| Aging Dam | 186 | 254782 | 186 | 633 | 2150 |
4.2 Baselines
We select three types of methods for comparison: 1) sparse retriever BM25 Robertson and Walker (1994); 2) dense embedding retrieval methods, Contriever Izacard et al. (2022), e5 Wang et al. (2024a), Nvidia/NV-Embedv2 Lee et al. (2025); 3) graph-based methods, GraphRAG Edge et al. (2024), LightRAG Guo et al. (2024), and HippoRAG Gutiérrez et al. (2024) and HippoRAG 2 Gutiérrez et al. (2025). All experiments were conducted on an A100 GPU with 80 GB memory. More details are presented in Appendix D.
| Method | Hurricane | Geography | Aging Dam | |||||||||
| Automatic Metric | LLM-as-Judge | Automatic Metric | LLM-as-Judge | Automatic Metric | LLM-as-Judge | |||||||
| BLEU | BERTScore | Correct | Complete | BLEU | BERTScore | Correct | Complete | BLEU | BERTScore | Correct | Complete | |
| No Retrieval | ||||||||||||
| GPT-4o | 2.6 | 66.8 | 57.1 | 23.5 | 2.8 | 63.9 | 64.2 | 42.1 | 5.5 | 62.5 | 56.9 | 19.6 |
| DeepSeek-R1 | 2.2 | 69.2 | 53.4 | 28.5 | 1.3 | 61.7 | 57.9 | 45.7 | 8.7 | 71.6 | 35.4 | 18.8 |
| Qwen2.5-7B-Instruct | 1.2 | 46.9 | 23.5 | 6.9 | 1.7 | 47.1 | 32.1 | 12.6 | 1.9 | 43.4 | 15.5 | 6.5 |
| Llama-3.3-70B-Instruct | 2.0 | 60.2 | 39.8 | 14.6 | 1.6 | 51.3 | 49.3 | 21.2 | 2.4 | 46.7 | 39.7 | 17.1 |
| Llama-4-Scout-17B-16E-Instruct | 2.3 | 62.7 | 53.9 | 39.1 | 3.1 | 61.3 | 48.8 | 21.9 | 5.2 | 64.2 | 34.4 | 15.1 |
| Sparse and Dense Embedding Retrieval | ||||||||||||
| BM25 Robertson and Walker (1994) | 10.5 | 77.5 | 74.3 | 60.8 | 7.9 | 73.3 | 79.7 | 70.8 | 18.1 | 80.2 | 85.7 | 75.9 |
| Contriever Izacard et al. (2021) | 11.3 | 78.6 | 75.7 | 60.6 | 7.5 | 74.1 | 82.5 | 73.6 | 16.7 | 76.8 | 70.9 | 60.9 |
| e5 Wang et al. (2024a) | 11.9 | 79.1 | 78.3 | 63.9 | 7.1 | 73.9 | 81.2 | 71.2 | 19.7 | 79.8 | 88.9 | 75.5 |
| NV-Embed Lee et al. (2025) | 12.2 | 79.1 | 76.9 | 59.6 | 7.5 | 75.4 | 81.9 | 72.3 | 18.8 | 79.6 | 79.4 | 73.6 |
| Graph-based RAG | ||||||||||||
| LightRAG Guo et al. (2024) | 7.6 | 72.2 | 64.1 | 47.6 | 7.0 | 67.8 | 69.8 | 59.6 | 12.8 | 75.4 | 65.9 | 46.5 |
| GraphRAG Edge et al. (2024) | 6.5 | 72.8 | 68.8 | 50.9 | 6.9 | 69.4 | 70.6 | 63.8 | 11.8 | 72.4 | 72.5 | 58.6 |
| HippoRAG Gutiérrez et al. (2024) | 6.8 | 73.6 | 70.5 | 48.8 | 7.8 | 70.3 | 76.9 | 69.7 | 12.4 | 73.8 | 80.9 | 58.6 |
| HippoRAG 2 Gutiérrez et al. (2025) | 8.6 | 76.3 | 86.7 | 53.5 | 9.2 | 74.6 | 80.6 | 67.5 | 11.5 | 74.1 | 88.2 | 53.7 |
| Our Method | ||||||||||||
| Hypercube-RAG | 13.2 | 82.5 | 80.6 | 66.2 | 9.3 | 76.8 | 83.6 | 74.8 | 21.5 | 82.6 | 86.9 | 77.8 |
4.3 Evaluation Metrics
Following the work Cui et al. (2025), we evaluate the quality of LLMs’ responses using automated metrics derived from token-level representations, including BLEU Papineni et al. (2002) and BertScore Zhang* et al. (2020) to provide the standard assessment of response quality. Since the settings are open-text responses, we also employ LLM-as-a-Judge that assesses correctness and completeness. All LLM-based evaluations use GPT-4o as the judge. The prompts used are in Appendix A.
5 Results and Analysis
5.1 Accuracy
We compare our Hypercube-RAG with semantic RAG methods, graph-based RAG methods, and LLMs without retrieval. Table 2 reveals three key findings. First, Hypercube-RAG consistently outperforms other baselines across three datasets, demonstrating its effectiveness in enhancing the capabilities of LLMs for in-domain scientific question-answering. Compared to the second-best method, Hypercube-RAG improves the accuracy by measured as relative gains across three of four metrics used. Second, all RAG methods perform much better than direct inference of LLMs without retrieval, as expected, showing the effectiveness of incorporating external knowledge for in-domain QA tasks. Finally, we also evaluate Hypercube-RAG against a range of baseline methods across multiple LLMs developed by different companies. As illustrated in Figure 5, our method achieves superior performance across all models regardless of the underlying LLM architectures.
5.2 Efficiency
We study the response speed of diverse RAG methods over the corpus size and report the retrieval time for one query in Table 3. We break down our analysis as follows. Within corpus size . BM25 achieves the fastest retrieval among all evaluated methods without surprise. In contrast, semantic and graph-based RAGs incur significantly higher retrieval costs. This is particularly pronounced for graph-based methods due to the computational burden of search paths in large-scale graphs, posing scalability challenges. Notably, our Hypercube-RAG substantially reduces the retrieval time by one to two orders of magnitude compared to both semantic and graph-based methods, underscoring the effectiveness of the hypercube structure in optimizing retrieval efficiency. We attribute this to the retrieval with compact cube labels introduced in Section 3.3. Although it is slightly slower than BM25, the marginal increase in retrieval time is a worthwhile trade-off for significantly improved accuracy, as shown in Table 2. Beyond corpus size . To mimic the real-world scenario that the external knowledge database is usually large and noisy, we expand the hurricane corpus from 844 to 11,539 documents (around 14 times). The added documents are related to other topics, such as “Pollution” or “Ozone Holes”, introducing additional noise. Hypercube-RAG maintains the lowest retrieval time per query (see the last column in Table 3), highlighting the efficiency of the cube label-based retrieval mechanism. This is likely because most noisy documents result in empty or sparsely populated cube cells, as they lack corresponding in-domain fine-grained labels pre-defined in our hypercube. As a result, despite a significantly expanded corpus, the hypercube can efficiently bypass these irrelevant documents. In contrast, other baseline methods exhibit greater sensitivity to corpus size, resulting in increased retrieval times.
| Methods | |||||
| BM25 | 0.5 | 0.9 | 1.6 | 3.2 | 51.1 |
| e5 | 14.4 | 17.4 | 19.1 | 37.8 | 82.4 |
| GraphRAG | 114.7 | 348.2 | 1179.4 | 5260.6 | 45135.7 |
| Hypercube-RAG | 0.7 | 1.5 | 3.3 | 7.1 | 22.1 |
In addition, Table 4 shows that the accuracy of the evaluated methods drops due to the inclusion of these noisy off-topic documents. However, Hypercube-RAG still outperforms others, demonstrating robust resilience to noisy data.
| Methods | BERTScore () | ||
| BM25 | 77.5 | 77.2 | 0.3 |
| e5 | 79.1 | 76.3 | 2.8 |
| GraphRAG | 72.8 | 70.5 | 2.3 |
| Hypercube-RAG | 82.5 | 80.3 | 1.9 |
5.3 Explainability
The retrieval process with our hypercube is inherently explainable through its associated cube labels. Considering the query “How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay?”, illustrated in Figure 4, Hypercube-RAG returns relevant documents by matching the decomposed components with fine-grained labels, ‘Melbourne Beach’, ‘Tropical Storm Fay’, and ‘rainfall’, along hypercube dimensions: LOCATION, EVENT, and THEME. Table 5 shows that three retrieved documents (Doc. 565, 246, 535), which are represented with fine-grained cube labels. These cube labels clearly explain why those documents are returned. For example, Doc. 565 is ranked the first to return, since it contains all three key components of the query. The detailed document contents are included in Appendix G.
| Doc. | Location | Event | Theme |
| 565 | ‘Melbourne Beach’: 1 | ‘Tropical Storm Fay’: 1 | ‘Rain’: 5 |
| 246 | ‘Florida’: 1 | ‘Tropical Storm Fay’: 1 | – |
| 535 | ‘Florida’: 1 | – | – |
5.4 Ablation Study
As shown in Section 3.3, Hypercube-RAG combines two retrieval strategies: sparse exact matching and dense embedding retrieval. Section 3.4 introduces the document ranking algorithm used to filter and prioritize the retrieved results. To study their effectiveness, we conduct an ablation study by removing each of them from Hypercube-RAG, which are represented as No-Sparse, No-Dense, and No-Ranking, respectively. Figure 7 shows that the full version of Hypercube-RAG consistently outperforms all other variants, verifying positive contributions of each constituent. More specifically, it justifies the benefit of coupling these two strategies for in-domain theme-specific question-answering tasks, and the necessity of the ranking algorithm to filter possible “noisy” documents. The ablation study on hypercube dimensions is in Appendix F.
5.5 Parameter Study
Hypercube-RAG combines sparse exact match and semantic retrieval strategies, as introduced in Section 3.3. We investigate the effect of the semantic retrieval strategy on answer quality under varying similarity thresholds, denoted by . Figure 8 shows that performance improves with increasing threshold values when , suggesting that higher similarity scores are more effective for retrieving semantically relevant documents. However, when , the behavior of semantic RAG begins to gradually degrade to the sparse exact matching strategy, resulting in a decline in performance. This is primarily due to overly strict matching criteria, which reduce the number of retrieved documents and thus limit the available context.
5.6 Case Study
An additional case study in Figure 6 presents that both Semantic-RAG and HippoRAG-2 fail to generate the correct answer due to inaccurate document retrieval. These methods tend to prioritize returning documents that are semantically similar in the embedding space. While the retrieved documents mention general topics such as above-normal, near-normal, and below-normal hurricane season, they miss the specific location ‘Atlantic’ and organization ‘Climate Prediction Center’. In contrast, our Hypercube-RAG returns the correct answer. It accurately identifies and retrieves documents containing key elements of the query, including ‘Atlantic’, ‘hurricane season’, and ‘Climate Prediction Center’, corresponding to the hypercube dimensions, LOCATON, THEME, and ORGANIZATION, respectively. This retrieval approach intuitively prioritizes specific contextual information over abstract, ambiguous references such as above-normal, near-normal, and below-normal alone.
6 Conclusions
In this work, we start by analyzing the strengths and limitations of existing RAG methods for in-domain theme-specific applications. To retain their advantages while addressing key challenges, we propose Hypercube-RAG, innovating retrieval based on fine-grained labels with a multidimensional structure (i.e., hypercube). The label-based retrieval integrates sparse lexical and dense semantic strategies, making it accurate, efficient, and inherently explainable. Empirical results demonstrate that our method consistently outperforms existing RAG baselines in accuracy and efficiency, while providing interpretable provenance.
Limitations
The current hypercube construction partially relies on manual input from domain experts to predefine different dimensions. To enhance generalizability across diverse scientific question-answering tasks, we aim to dynamically design the hypercube dimensions. One possibility is to leverage large language models (LLMs) to automate this process.
Ethics Statement
We conduct our experiments on synthesized QA datasets collected from a publicly available scientific corpus. Our proposed approach and findings do not pose any foreseeable harm to individuals or groups. Overall, we do not anticipate any significant ethical concerns with this work.
Acknowledgments
This work is supported by the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE), which is funded by the National Science Foundation (NSF) under award number: 2118329. Any opinions, findings, conclusions, or recommendations expressed herein are those of the authors and do not necessarily represent the views of NSF.
Appendix
Appendix A Prompt Templates
Appendix B QA Categories
Appendix C Discussion
C.1 Why does Hypercube-RAG outperform other RAG methods?
Our Hypercube-RAG is particularly proposed for domain-focused analysis, where domain-specific information, e.g., location, date, event, event, theme, is central to the query. In such cases with theme-specific information, our hypercube enables precise and efficient retrieval by accessing only one or a few cells in a theme-specific hypercube, as each cell encapsulates multiple dimensions of information associated with a document. These fine-grained document labels can be represented in a hypercube along different dimensions. When a query comes in, accessing one or a few hypercubes presents a higher chance of retrieving highly relevant documents covering all key information contained in a query - Accuracy. Additionally, retrieval is operated with the cube labels significantly speeds up the retrieval time since those labels are in a compact space - Efficiency. In contrast, sparse exact match and dense embedding retrieval require traversing all documents and contents inside. Moreover, the retrieval process is inherently explainable, as the searching is based on the document’s labels represented by a hypercube - Explainability.
Case 1: simple query.
How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay? The relevant documents can be retrieved from one cube cell in a hypercube <Melbourne Beach, rainfall, Tropical Storm Fay >.
Case 2: long query with multiple topics.
In cases where a query is very diverse, they may need to access multiple cube cells such that the query information can be covered as much as possible. For example, the diverse query could be “What consequences were caused by Tropical Storm Fay, such as how much rainfall Melbourne Beach and Palm Bay in Florida received, and how much monetary losses were caused?”. Those cube cells needed could be <Melbourne Beach, rainfall, Tropical Storm Fay >, <Palm Beach, rainfall, Tropical Storm Fay >, <Melbourne Beach, monetary losses, Tropical Storm Fay >, <Palm Beach, monetary losses, Tropical Storm Fay>.
Case 3: diverse query with shifted topics.
Given a more complicated query including shifted topics, e.g., “How did the monetary loss caused by Tropical Storm Fay impact the industrial layoff in Miami Downtown?”, cube cells across two hypercubes are needed <Miami Downtown, Industrial Layoff, Tropical Storm Fay > and <Miami Downtown, Monetary loss, Tropical Storm Fay > to get more relevant documents.
C.2 Why did we use in-domain scientific QA datasets rather than general ones?
According to the above discussion, hypercube construction is sensitive to themes in various domains. While similar or related themes are allowed to be included in a hypercube, each theme-specific dataset should have a separate hypercube to maximize its power. However, the general QA datasets, e.g., PopQA Mallen et al. (2022) and NQ Kwiatkowski et al. (2019), typically present articles across multiple themes. The process of constructing multiple hypercubes will require intensive efforts (e.g., design hypercube schema for different datasets.), which is out of the scope of this study. This is why we particularly focus on in-domain theme-specific QA datasets.
Appendix D Implementation Details
We selected three types of RAG baseline methods for comparison: sparse retriever BM25, dense embedding retrieval, and graph-based methods.
For sparse retriever BM25, we implement it using a GitHub repository444https://github.com/dorianbrown/rank_bm25.
For most experiments, we used the GPT-4o555https://platform.openai.com/docs/models/gpt-4o as the LLM base unless explicitly specified. In Table 2, we also include baseline methods without RAG, such as DeepSeek-R1666https://github.com/deepseek-ai/DeepSeek-R1, Qwen2.5-7B-Instruct777https://qwenlm.github.io/blog/qwen2.5/, Llama-3.3-70B-Instruct888https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, and Llama-4-Scout-17B-16E-Instruct999https://huggingface.co/meta-llama/Llama-4-Scout-17B-16E-Instruct. Note that DeepSeek-R1 is a reasoning model, we set 0.6 and 1 as the temperature by following their recommendations. Since its output also includes the reasoning/thinking process, we exclude it and consider the final output as the answer. The temperatures for other models are set to 0.
In the evaluation phase, we compute BLEU scores and BERTScores with the NLTK101010https://www.nltk.org/ and SentenceTransformer111111https://sbert.net/ Python packages, respectively. We also use an LLM (GPT-4o) as the judge to compute the correctness and comprehensiveness (see the prompt in Appendix A).
Appendix E Details on Theme-specific Extraction
A small number of specialized terms are manually added, such as “hurricane landfall”, “hurricane hitting”, “hurricane intensity”, “sustained winds”.
Appendix F Ablation Study on Hypercube Dimensions
We conduct an ablation study on the hypercube dimensions: Location, Event, Date, Organization, Person, and Theme. To study their effectiveness, we conduct an ablation study by removing each of the dimensions, which are represented as No-Location, No-Event, No-Date, No-Organization, No-Person, and No-Theme, respectively. Table 6 shows that the full version of Hypercube-RAG consistently outperforms all other variants, verifying positive contributions of each dimension. More specifically, we observe that the performance drops the most when removing the Location and Theme dimensions, justifying the high importance of these two hypercube dimensions for in-domain theme-specific question-answering.
| Dimension | Hurricane | Geography | Aging Dam |
| No-Location | 78.7 | 65.9 | 77.5 |
| No-Event | 79.6 | 75.9 | 80.0 |
| No-Date | 79.7 | 74.5 | 80.4 |
| No-Organization | 79.6 | 74.6 | 80.2 |
| No-Person | 80.2 | 73.2 | 81.4 |
| No-Theme | 77.9 | 71.7 | 81.1 |
| FULL | 82.5 | 76.8 | 82.6 |
Appendix G Document Contents
| Document ID | Document Content |
| 565 | …Why is September the peak month for hurricanes? NASA oceanographer, Bill Patzert at the Jet Propulsion Laboratory, Pasadena, Calif., provided the answer: “Hurricanes are fueled by warm ocean temperatures and September is the end of the Northern Hemisphere ocean warming season. The 2008 Atlantic hurricane season started early with the formation of Tropical Storm Arthur on May 30, from the remnants of the eastern Pacific Ocean’s first storm, Alma, which crossed Central America and reformed in the Gulf of Mexico. It took one month and four days for the next storm to form, Bertha. Now, the action is definitely picking up. The tropical Atlantic is warm, but not unusually so.” Once a powerful Category 3 hurricane, now a tropical depression, Gustav moved from northwest Louisiana into northeastern Texas and into Arkansas by Sept. 3. Like Tropical Storm Fay in August, Gustav’s legacy will lie in large rainfall totals. According to the National Hurricane Center discussion on Sept. 2, "Storm total rainfalls are expected to be five to ten Inches with isolated maximums of 15 inches over portions of Louisiana, Arkansas and Mississippi. Rainfall amounts of 4-8 inches have been already reported in parts of Alabama, Mississippi and Louisiana. …On Sept. 4, Ike strengthened into a major hurricane, Category 4 on the Saffir-Simpson Scale, with maximum sustained winds near 145 mph. Ike is forecast to head west and may also affect the Bahamas. Behind Ike, on Sept. 2, the tenth tropical depression in the Atlantic Ocean basin was born. Tropical Depression 10 formed west of the African coast, so it has a long way to go before it has any impact on the U.S. or the Caribbean. By the late morning, that tropical cyclone became Tropical Storm Josephine. In August, Fay’s ten-day romp from the U.S. Southeast northward up the Appalachian Mountains seemed like a harbinger for September’s storms. Fay took her time going northward and dumped tremendous amounts of rain along the way. Melbourne Beach, Florida, received as much as 25.28 inches of rain. Other cities in various states reported high totals: Thomasville, Ga., reported 17.43 inches; Camden, Ala., received 6.85 inches; Beaufort, S.C., received 6.11 inches; Carthage, Tenn., reported 5.30 inches, and Charlotte, N.C., reported 5.90 inches. Fay was a perfect example of how weaker tropical storms can cause flooding inland … |
| 246 | …Among the already apparent evidence: Dunes that historically protected Kennedy Space Center from high seas even during the worst storms were leveled during Tropical Storm Fay in 2008, Hurricane Irene in 2011 and Hurricane Sandy in 2012. A stretch of beachfront railroad track built by NASA in the early 1960s that runs parallel to the shoreline has been topped by waves repeatedly during recent storms. …The research came about after NASA partnered with the U.S. Geological Survey and UF to figure out why chronic erosion was happening along a roughly 6-mile stretch of beach between launch pads 39A and 39B – the ones used for Space Shuttle and Apollo missions. The problem had been occurring for years but seemed to be growing worse, beginning with the spate of hurricanes that struck Florida in 2004. Jaeger said he, Adams and doctoral students Shaun Kline and Rich Mackenzie determined the cause was a gap in a near-shore sandbar that funnels the sea toward that section of beach. … |
| 535 | …The topic has local interest because it has recently been observed that significant erosion is occurring on New Orleans area levees primarily caused by seepage-driven flow. Recently work began by the U.S. Army Corps of Engineers to plug in the shipping channel known as the Mississippi River-Gulf Outlet Canal (MRGO) that extends for 76 miles from New Orleans through wetlands to the Gulf of Mexico. The erosion of thousands of acres of cypress wetlands and marsh caused by the channel has been blamed for increased storm surge during Hurricane Katrina in 2005. Our theory would suggest that seepage caused by underwater flow will continue to erode and weaken the levee system around New Orleans, but the rate of this erosion should gradually slow with time, says Straub. Hopefully this research will aid the U.S. Army Corps of Engineers in identifying levees that need repair and assessing the lifespan of structures like the MRGO that are not planned for upkeep. Using fieldwork conducted in the Florida Panhandle, Straub and his fellow researchers were able to better understand the process of seepage erosion, which occurs when the re-emergence of groundwater at the surface shapes the Earth’s topography. The paper suggests that the velocity at which channel heads advance is proportional to the flux of groundwater to the heads. The researchers used field observations and numerical modeling to come up with the theory. … |
Appendix H AI Assistants Usage
We achieve most work by ourselves while using ChatGPT to partially provide coding help (e.g., grammar checks, package searches).