### v2

# Hypercube-Based Retrieval-Augmented Generation for Scientific Question-Answering

Abstract.
Large language models (LLMs) often need to incorporate external knowledge to solve theme-specific problems. Retrieval-augmented generation (RAG) has shown its high promise, empowering LLMs to generate more qualified responses with retrieved external data and knowledge. However, most RAG methods retrieve relevant documents based on either sparse or dense retrieval methods or their combinations, which overlooks the essential, multi-dimensional, and structured semantic information present in documents. This structured information plays a critical role in finding concise yet highly relevant information for domain knowledge-intensive tasks, such as scientific question-answering (QA). In this work, we introduce a multi-dimensional (cube) structure, Hypercube, which can index and allocate documents in a pre-defined multi-dimensional space. Built on the hypercube, we further propose Hypercube-RAG, a novel RAG framework for precise and efficient retrieval. Given a query, Hypercube-RAG first decomposes it based on its entities, phrases, and topics along with pre-defined hypercube dimensions, and then retrieves relevant documents from cubes by aligning these decomposed components with corresponding dimensions. Experiments on three datasets across different domains demonstrate that our method improves response accuracy by and retrieval accuracy by over the strongest RAG baseline. It also boosts retrieval efficiency (speed) by one or two magnitudes faster than graph-based RAG. Notably, our Hypercube-RAG inherently offers explainability by revealing those underlying dimensions used for retrieval. The data and code are available at https://github.com/JimengShi/Hypercube-RAG.
1. Introduction
Hypecube- vs. Semantic RAG: A case study
Large language models (LLMs) often suffer from hallucinations and factual inaccuracies, especially for scientific question-answering tasks (Augenstein et al., 2024). To address this phenomenon, retrieval-augmented generation (RAG) has emerged as the de facto approach, incorporating external domain knowledge to generate contextually relevant responses (Huang et al., 2025). Despite recent advancements, they exhibit key limitations in theme-specific applications on which the fine-grained topics are highly focused. For instance, retrieved documents are often semantically similar to the query but overlook specialized themes (e.g., terminology or nuanced contextual cues) prevalent in specific literature or reports. Moreover, the retrieval process often suffers from inefficiency and limited transparency, which are critical in high-stakes domains, such as healthcare (Pujari et al., 2023), chemistry (Wellawatte et al., 2025), environment (Wang et al., 2025), and law (Pipitone and Alami, 2024).
The conventional sparse lexical retrieval ranks a set of documents based on the appearance of query terms (Kadhim, 2019). Due to its reliance on the exact token overlap, it precisely captures specific themes or topics, and the retrieval process is efficient and interpretable. However, it struggles to retrieve contextually related or paraphrased documents that require semantic understanding. To mitigate the above limitation, dense embedding retrievers were proposed by computing the similarity between vector embeddings of query-document pairs (Karpukhin et al., 2020). While contextual understanding is relatively improved, the hurdle of missing specific themes remains, causing semantically similar but off-topic retrievals (Krishna et al., 2021; Kang et al., 2024). Figure 1 shows an example where the Semantic-RAG returns documents related to (or the lack of) precipitation in Florida, but overlooks the critical location-specific information for “Melbourne Beach”. Moreover, retrieval in an embedding space makes it challenging to interpret why certain documents were selected (Ji et al., 2019). Additionally, a graph-based RAG represents documents as graph structures, where nodes correspond to entities or concepts and edges capture their relationships (Peng et al., 2024). They offer explanations of the retrieval process by outputting the traversed subgraph. However, two limitations persist still. First, traversing such subgraphs introduces a significant risk of information overload, as the expanding neighborhoods may include substantial amounts of irrelevant themes (Wang et al., 2024a). Second, graph-based retrieval methods are prone to inefficiencies and considerable scalability bottlenecks (Devezas and Nunes, 2021; Zhang et al., 2025a; Han et al., 2024), primarily stemming from the vast number of nodes and edges in their underlying graph structures.
The methods discussed above tend to perform poorly in at least one of the aspects among accuracy, efficiency, and explainability. This motivates us to develop a RAG system that overcomes this weakness. In this pursuit, we identified the text cube as a promising technique (Tao et al., 2018; Wang et al., 2023). A text cube is an inherently explainable multidimensional structure that allocates documents into cubes along various pre-defined dimensions, such as location, date, event, and specific-theme(s). Each cell is associated with fine-grained entities (e.g., “Miami Downtown”, “2021”, “Hurricane Irma”, “flooding”) extracted from documents, which are viewed as document labels. Relevant documents can be indexed and located in cube cells associated with these document labels (cube labels for brevity). Moreover, the theme dimension specifies the themes in hierarchies which make the text cube well-suited for theme-specific applications where users primarily focus on fine-grained inquiries (Tao et al., 2018), such as which hurricane hits and how much rainfall it results in certain areas, or which medicine is effective in treating which type of cancer.
In this work, we implement a multi-dimensional (cube) structure called Hypercube, which indexes documents to cube cells based on their fine-grained labels in a pre-defined, multi-dimensional semantic space. Building on this cube structure, we further propose Hypercube-RAG, a novel RAG framework to enhance the retrieval process by retrieving relevant documents from the corresponding cube cells. Compared with the aforementioned RAG approaches, our Hypercube-RAG simultaneously exhibits three critical characteristics due to its multidimensionality: (1) High Accuracy. The multidimensional, cube label-based search supports both sparse and dense embedding strategies, effective for capturing essential information in multi-dimensional space as well as uncommon thematic terminology and semantically relevant cases. (2) High Efficiency. The retrieval mechanism quickly narrows down the search to the pertinent cube cell(s) based on their associated labels. This operation exhibits a constant time complexity of , which could be multiple magnitudes faster, thereby demonstrating potential robust scalability concerning corpus size. (3) Explainability. Hypercube is inherently explainable as fine-grained labels associated in cube cells represent the compact information in documents (see Figure 2 (right) and Figure 4). Overall, our contributions are as follows:
We identify the shortcomings of conventional RAG approaches, especially for theme-specific question-answering tasks in the scientific domains.
We design a multi-dimensional structure (hypercube) and further propose Hypercube-RAG, a simple yet accurate, efficient, and explainable RAG method.
We conduct experiments on three datasets across different domains with multiple metrics, demonstrating that our approach outperforms other baseline methods consistently.
2. Related Work
2.1. Text-based RAG
BM25 (Robertson and Walker, 1994; Trotman et al., 2014) is a classical sparse retrieval method based on TF-IDF principles, widely used in information retrieval. It ranks documents by scoring query term matches, with adjustments for term frequency saturation and document length. While efficient and interpretable, BM25 lacks semantic understanding and often misses paraphrased or contextually related documents without exact lexical overlap, limiting its effectiveness in tasks requiring deeper language comprehension. Dense embedding retrievers compute contextual similarity between the given query and a document/chunk in the embedding space. Common retrievers consist of DPR (Karpukhin et al., 2020), Contriever (Izacard et al., 2022), e5 (Wang et al., 2022), and ANCE (Xiong et al., 2021). By capturing deeper semantic relationships, these methods improve performance in open-domain question answering and generative tasks, especially when surface-form matching fails. However, dense retrieval can return semantically close yet topically irrelevant documents, introducing noise and hallucinations. Moreover, its opaque nature also hinders interpretability (Ji et al., 2019), limiting adoption in high-stakes or scientific domains. Recent efforts aim to mitigate these limitations by incorporating hybrid retrieval techniques (Maillard et al., 2021) and domain-adaptive retrievers (Lee et al., 2021).
2.2. Structured RAG
Graph-based RAG methods enhance the standard RAG paradigm by introducing structured knowledge representations, such as entity or document graphs. GraphRAG (Edge et al., 2024) derives an entity knowledge graph from the source documents and gathers summaries for all groups of closely related entities. LightRAG (Guo et al., 2024) employs a dual-level retrieval system that enhances comprehensive information retrieval from both low-level and high-level knowledge discovery. HippoRAG (Gutiérrez et al., 2024) is a knowledge graph-based retrieval framework inspired by the hippocampal indexing theory of human long-term memory. HippoRAG 2 (Gutiérrez et al., 2025) builds on top of HippoRAG and enhances it with deeper passage integration and more effective online use of an LLM. SELF-RAG (Asai et al., 2023) enhances an LLM’s quality and factuality through retrieval and self-reflection. RA-DIT (Lin et al., 2023) fine-tunes the pre-trained LLMs or retrievers to incorporate more up-to-date and relevant knowledge. Despite these improvements, graph construction and reasoning can be computationally intensive due to the vast number of nodes and edges.
Text Cube. The text cube has been proposed to structure spatial-related data across geospatial dimensions, thematic categories, and diverse application semantics (Wang et al., 2023). STREAMCUBE implements the text cube to incorporate both spatial and temporal hierarchies (Feng et al., 2015). Doc2Cube (Tao et al., 2018) automates the allocation of documents into a text cube to support multidimensional text analytics. Despite these elaborate designs, the text cube has not yet been explored in the context of retrieval-augmented generation (RAG) systems.
3. Our Method: Hypercube-RAG
Figure 3 illustrates the Hypercube-RAG framework through a concrete question-answering example. In the following, we define the hypercube structure (Sec. 3.1), describe the process for designing and constructing the hypercube (Sec. 3.2), explain how two sparse and dense embedding retrieval strategies are combined inside of hypercube (Sec. 3.3), and present how to rank retrieved documents (Sec. 3.4).
3.1. Hypercube Formulation
Given a text corpus with documents , the hypercube is designed as a multidimensional structure, , where is the number of hypercube dimensions, and is the dimension. For any document , a text classification algorithm (e.g, TeleClass (Zhang et al., 2025b)) will assign it to one or more cube cells. This process is equivalent to assigning a -dimensional labels () for a document, where label represents a category of document in dimension .
3.2. Hypercube Design and Construction
Hypercube Design. The core idea of the hypercube is to identify fine-grained entities or topics present in the corpus and to index documents into cube cells along appropriate dimensions, thereby enabling the subsequent retrieval. In this work, we design the hypercube structure by systematically determining its dimensions. First, we employ pre-trained language models to extract entities from the entire corpus. Next, the K-means clustering algorithm is applied to group these entities into semantically coherent clusters. For each cluster, we let a large language model (LLM) generate concise summaries that characterize the underlying cluster categories or dimensions111For example, LLM outputted “these entities in the cluster include countries, states, provinces, cities, towns, and specific addresses.”. Finally, we consolidate the candidate dimensions across clusters to construct a high-level, multi-dimensional space for document indexing.
This dimension-identification process for hypercube design has demonstrated strong alignment with domain expert expectations and yielded promising performance (although diverse strategies remain open for future exploration).
In Appendix A, we provide a detailed dimension-identification process for each dataset.
Hypercube Construction. With the hypercube dimensions finalized or pre-defined, we refine the previously generated concise summaries as prompts to an LLM to extract entities associated with each high-level dimension. These extracted entities are then used as labels to index documents into their corresponding cube cells. The process to extract entities for one dimension is represented as:
where represents the extracted entity set from a document, , along the dimension, , refers to an LLM, and denotes the corresponding prompt used for . We present a prompt example in Appendix B.
3.3. Hypercube Retrieval
A constructed knowledge hypercube can serve for the retrieval process based on its fine-grained labels. Given a query , an LLM first decomposes it into a set of entities, , aligned with pre-defined dimensions (see the prompt in Figure 3), then retrieval inside of hypercube is performed by matching these decomposed components with labels associated in cube cell(s). Our hypercube supports two matching strategies: sparse exact lexical match and dense embedding match. We prioritize leveraging the exact matching strategy due to its high precision, scoring documents based on the frequency with which query entities appear:
where is an indicator function that equals 1 if an entity of the query exactly matches a term in the document, and 0 otherwise.
Considering the example in Figure 3, the last two components from the query, Melbourne Beach and Tropical Storm Fay, are handled using the exact matching strategy, as they correspond precisely to fine-grained labels in the hypercube. However, the first component, Rainfall, does not exactly match any label(s) in the hypercube, even though its matched label should be Rain.
To address such cases, Hypercube-RAG supports semantic retrieval by computing the similarity between query components and fine-grained cube labels, along with a certain dimension, . When the similarity score exceeds a predefined threshold , semantic retrieval is triggered. Specifically, both the query entities and cube labels are projected into an embedding space, , with an encoder. It can be represented as:
where refers to the similarity function.
3.4. Hypercube Ranking
The retrieved documents need to be provided to the LLM as contextual input. However, long-context RAG methods do not always improve the quality of LLM responses, since they may introduce irrelevant or noisy information (Liu et al., 2024; Jin et al., 2025). This raises a critical challenge: how to precisely filter and select theme-relevant documents. To address this, we prioritize returning those documents that fully cover key components derived from the query:
If no documents fully cover all query components (i.e., ), we return the next best set of documents with the highest partial coverage:
4. Experiments
4.1. Datasets
To verify the applicability of Hypercube-RAG, three datasets over varying types across different domains are selected: (1) SciFact (Wadden et al., 2020): a dataset of expert-written scientific claims paired with evidence-containing biomedical abstracts annotated with labels and rationales. (2) LegalBench (Pipitone and Alami, 2024): a dedicated QA and retrieval benchmark in the legal domain, entirely human-annotated by legal experts. (3) SciDCC (Mishra and Mittal, 2021): a synthesized dataset using LLMs in the geo-science domain, covering “Geography”, “Hurricane”, and “aging dams”. We collected the text corpus from a subset of the Science Daily Climate Change dataset (SciDCC) (Mishra and Mittal, 2021) from the Science Daily website222https://www.sciencedaily.com/. “Aging dams” discusses dam failures in the United States, whose corpus comprises related news articles downloaded from Google News333https://news.google.com/, collected by domain experts. The summary is presented in Table 1.
| Datasets | # Docs | # QA | QA type | Ques. Length | Ans. Length |
| SciFact (Medical) | 5813 | 188 | True or False | 528 | 1 |
| LegalBench (Law) | 95 | 977 | Open-Ended | 1639 | 10304 |
| SciDCC (Geoscience) | 1462 | 789 | Open-Ended | 636 | 10150 |
In this work, we focus on studying Hypercube-RAG with one hypercube for each dataset, while more complex domains across various themes may benefit from multiple hypercubes. We provide a detailed discussion on multiple hypercubes in Appendix C.
| Method | SciFact (Medical) | LegalBench (Law) | SciDCC (Geoscience) | |||||||||
[…]
| F1 | Semantic | Correctness | Completeness | F1 | Semantic | Correctness | Completeness | F1 | Semantic | Correctness | Completeness | |
[…]
| GPT-4o | 81.3 | 84.9 | 81.9 | 81.9 | 22.6 | 56.8 | 52.7 | 32.1 | 26.5 | 64.6 | 59.8 | 29.7 |
[…]
| BM25 (Robertson and Walker, 1994) | 87.8 | 89.8 | 87.7 | 87.7 | 31.8 | 61.2 | 54.6 | 50.9 | 45.3 | 76.6 | 79.1 | 68.2 |
| Contriever (Izacard et al., 2021) | 72.3 | 77.1 | 72.3 | 72.3 | 37.8 | 62.4 | 67.1 | 64.3 | 42.6 | 76.5 | 77.2 | 65.6 |
| e5 (Wang et al., 2024b) | 90.4 | 92.1 | 90.4 | 90.4 | 38.3 | 63.2 | 74.1 | 71.5 | 45.0 | 77.3 | 81.9 | 69.4 |
| NV-Embed-v2 (Lee et al., 2025) | 91.2 | 92.6 | 91.3 | 91.3 | 32.8 | 60.7 | 54.1 | 50.3 | 44.9 | 77.8 | 80.4 | 69.8 |
[…]
| LightRAG (Guo et al., 2024) | 67.3 | 69.3 | 76.6 | 76.6 | 28.5 | 61.2 | 68.9 | 66.5 | 30.9 | 68.3 | 66.7 | 51.9 |
| GraphRAG (Edge et al., 2024) | 84.9 | 86.2 | 84.8 | 84.8 | 36.3 | 61.8 | 68.4 | 64.7 | 39.2 | 71.4 | 70.4 | 56.6 |
| HippoRAG (Gutiérrez et al., 2024) | 85.2 | 87.3 | 86.1 | 86.1 | 34.7 | 60.4 | 67.5 | 63.5 | 41.6 | 72.4 | 75.4 | 59.1 |
| HippoRAG 2 (Gutiérrez et al., 2025) | 88.3 | 90.2 | 89.3 | 89.3 | 37.2 | 62.6 | 70.9 | 69.3 | 43.1 | 75.1 | 84.7 | 58.9 |
[…]
| Hypercube-RAG | 91.5 | 93.0 | 91.5 | 91.5 | 40.5 | 64.6 | 81.5 | 78.1 | 46.4 | 80.3 | 83.3 | 72.2 |
| Method | SciFact (Medical) | LegalBench (Law) | SciDCC (Geoscience) | |||||||||
[…]
| Precision@3 | Recall@3 | Precision@5 | Recall@5 | Precision@3 | Recall@3 | Precision@5 | Recall@5 | Precision@3 | Recall@3 | Precision@5 | Recall@5 | |
[…]
| BM25 (Robertson and Walker, 1994) | 27.5 | 75.9 | 18.1 | 81.3 | 12.0 | 36.1 | 7.8 | 38.8 | 28.6 | 86.0 | 17.9 | 89.7 |
| Contriever (Izacard et al., 2021) | 13.8 | 35.9 | 8.9 | 38.1 | 19.2 | 57.5 | 12.7 | 63.3 | 17.8 | 53.2 | 12.7 | 62.4 |
| e5 (Wang et al., 2024b) | 33.0 | 87.7 | 21.2 | 92.8 | 21.7 | 65.1 | 13.9 | 69.5 | 29.0 | 87.1 | 18.4 | 91.2 |
| NV-Embed-v2 (Lee et al., 2025) | 34.4 | 87.9 | 22.9 | 95.2 | 13.4 | 40.3 | 8.7 | 43.6 | 19.7 | 59.4 | 12.5 | 62.4 |
| HippoRAG 2 (Gutiérrez et al., 2025) | 34.7 | 88.2 | 22.6 | 95.6 | 12.8 | 38.8 | 12.7 | 64.3 | 18.2 | 57.6 | 12.1 | 61.7 |
[…]
| Hypercube-RAG | 37.5 | 88.8 | 27.6 | 96.1 | 26.1 | 75.4 | 16.9 | 79.7 | 35.5 | 86.0 | 28.8 | 92.4 |
4.2. Baselines
We select four types of methods for comparison: 1) sparse retriever BM25 (Robertson and Walker, 1994); 2) dense embedding retrieval methods: Contriever (Izacard et al., 2022), e5 (Wang et al., 2024b), Nvidia/NV-Embedv2 (Lee et al., 2025); 3) graph-based methods: GraphRAG (Edge et al., 2024), LightRAG (Guo et al., 2024), and HippoRAG (Gutiérrez et al., 2024) and HippoRAG 2 (Gutiérrez et al., 2025); and (4) LLMs without retrieval. We implement the sparse BM25 retriever using a GitHub repository444https://github.com/dorianbrown/rank_bm25. For dense embedding retrievers, we utilize their models uploaded to Hugging Face. Experiments for graph-based baselines using their official GitHub repositories. All experiments were conducted on a single A100 GPU with 80 GB memory.
4.3. Evaluation Metrics
We utilize multiple metrics for a comprehensive evaluation of different methods. To measure the QA performance, following the work (Cui et al., 2025), we evaluate the quality of LLMs’ responses using automated metrics derived from token-level representations, including F1 and Semantic score (Chandrasekaran and Mago, 2021) to provide the standard assessment of response quality. Since the settings are open-text responses, we also employ LLM-as-a-Judge that assesses correctness and completeness. All LLM-based evaluations use GPT-4o as the judge. The prompts and other parameters used are in Appendix B. To measure the QA performance, following the HippoRAG work (Gutiérrez et al., 2025), precison@ and recall@ rates are computed to measure the retrieval performance.
5. Results
5.1. Accuracy
5.1.1. QA performance.
Table 4.1 reports the results for QA performance. The observations and findings are as follows. We noticed that the state-of-the-art baselines present varying performance over different datasets. Most of them perform better on SciFact than the other two datasets. Notably, Hypercube-RAG consistently outperforms other baselines across all three datasets, demonstrating its effectiveness in enhancing the capabilities of LLMs for scientific question-answering tasks. Hypercube-RAG surpasses the second-best models with an average improvement by on F1 score, on semantic score, on correctness score, and on completeness score, evaluated across three datasets. The most significant improvements were observed on LLM-as-judge scores over the LegalBench and SciDCC datasets. We also observed that all RAG methods perform much better than direct inference of LLMs without retrieval, as expected, showing the effectiveness of incorporating external knowledge for scientific QA tasks.
5.1.2. Retrieval performance.
Table 4.1 reports retrieval results for datasets with supporting passage annotations and models that explicitly retrieve passages. The retrieval accuracy of baseline methods also varied by dataset: HippoRAG 2 was the second-best model on SciFact, while e5 achieved this rank on both LegalBench and SciDCC. Compared to the strongest baselines, Hypercube-RAG exhibits superior precision and recall scores with an average improvement by and on SciFact, and on LegalBench, and and on SciDCC, respectively. These results highlight Hypercube-RAG as a state-of-the-art RAG system that enhances both retrieval and QA performance.
5.2. Efficiency
We study the response speed of diverse RAG methods over the corpus size and report the retrieval time for one query in Table 4.
5.2.1. Within corpus size .
BM25 achieves the fastest retrieval among all evaluated methods without surprise. In contrast, semantic and graph-based RAGs incur significantly higher retrieval costs. This is particularly pronounced for graph-based methods due to the computational burden of search paths in large-scale graph structures, posing scalability challenges. Notably, our Hypercube-RAG substantially reduces the retrieval time by one to two orders of magnitude compared to both semantic and graph-based methods, underscoring the effectiveness of the hypercube structure in optimizing retrieval efficiency. We attribute this to the retrieval with compact, multi-dimensional cube labels introduced in Section 3.3. Although it is slightly slower than BM25, the marginal increase in retrieval time is a worthwhile trade-off for significantly improved retrieval accuracy and response quality, as shown in Tables 4.1 and 4.1.
| Methods | |||||
| BM25 | 0.5 | 0.9 | 1.6 | 3.2 | 51.1 |
| e5 | 14.4 | 17.4 | 19.1 | 37.8 | 82.4 |
| GraphRAG | 114.7 | 348.2 | 1179.4 | 5260.6 | 45135.7 |
| HippoRAG 2 | 1575.0† | 850.0 | 1722.8 | 2403.6 | 7490.4 |
| Hypercube-RAG | 0.7 | 1.5 | 3.3 | 7.1 | 22.1 |
5.2.2. Beyond corpus size .
To mimic the real-world scenario that the external knowledge database is usually large and noisy, we expand the hurricane-related corpus from 844 to 11,539 documents (around 14 times larger). The added documents are related to other topics, such as “Pollution” or “Ozone Holes”, introducing additional irrelevant documents (“noise”). Hypercube-RAG maintains the lowest retrieval time per query (see the last column in Table 4), highlighting the efficiency of the cube label-based retrieval mechanism. This is likely because most noisy documents result in empty or sparsely populated cube cells, as they lack corresponding fine-grained labels pre-defined in our hypercube. As a result, despite a significantly expanded corpus, the hypercube can efficiently bypass these irrelevant documents. In contrast, other baseline methods exhibit greater sensitivity to corpus size, resulting in increased retrieval times. In addition, Table 5 shows that the accuracy of the evaluated methods drops due to the inclusion of these noisy off-topic documents. However, Hypercube-RAG still outperforms others, demonstrating robust resilience to noise.
| Methods | Semantic Score with | Semantic Score with | Score () |
| BM25 | 77.5 | 77.2 | 0.3 |
| e5 | 79.1 | 76.3 | 2.8 |
| GraphRAG | 72.8 | 70.5 | 2.3 |
| HippoRAG 2 | 76.3 | 73.7 | 2.6 |
| Hypercube-RAG | 82.5 | 80.3 | 1.9 |
5.3. Explainability
The retrieval process with our hypercube is inherently explainable through its associated labels in the cube cell(s). Considering the query “How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay?” in Figure 3, Hypercube-RAG returns relevant documents by matching the decomposed components with fine-grained labels, “Melbourne Beach”, “Tropical Storm Fay”, and “rainfall”, along hypercube dimensions: Location, Event, and Theme. Table 6 shows that three retrieved documents (Doc. 565, 246, 535), which are represented with fine-grained cube labels. These cube labels clearly explain why those documents are returned. For example, Doc. 565 is ranked the first to return, since it contains all three key components of the query. The detailed document contents are included in Table 7.
| Doc. | Location | Event | Theme |
| 565 | ‘Melbourne Beach’: 1 | ‘Tropical Storm Fay’: 1 | ‘Rain’: 5 |
| 246 | ‘Florida’: 1 | ‘Tropical Storm Fay’: 1 | – |
| 535 | ‘Florida’: 1 | – | – |
| Document ID | Document Content |
| 565 | …The 2008 Atlantic hurricane season started early with the formation of Tropical Storm Arthur on May 30, from the remnants of the eastern Pacific Ocean’s first storm, Alma, which crossed Central America and reformed in the Gulf of Mexico. It took one month and four days for the next storm to form, Bertha. Once a powerful Category 3 hurricane, now a tropical depression, Gustav moved from northwest Louisiana into northeastern Texas and into Arkansas by Sept. 3. Like Tropical Storm Fay in August, Gustav’s legacy will lie in large rainfall totals. According to the National Hurricane Center discussion on Sept. 2, ”Storm total rainfalls are expected to be five to ten Inches with isolated maximums of 15 inches over portions of Louisiana, Arkansas and Mississippi. Rainfall amounts of 4-8 inches have been already reported in parts of Alabama, Mississippi and Louisiana. …By the late morning, that tropical cyclone became Tropical Storm Josephine. In August, Fay’s ten-day romp from the U.S. Southeast northward up the Appalachian Mountains seemed like a harbinger for September’s storms. Fay took her time going northward and dumped tremendous amounts of rain along the way. Melbourne Beach, Florida, received as much as 25.28 inches of rain. Other cities in various states reported high totals: Thomasville, Ga., reported 17.43 inches; Camden, Ala., received 6.85 inches; Beaufort, S.C., received 6.11 inches; Carthage, Tenn., reported 5.30 inches, and Charlotte, N.C., reported 5.90 inches. Fay was a perfect example of how weaker tropical storms can cause flooding inland … |
| 246 | …Among the already apparent evidence: Dunes that historically protected Kennedy Space Center from high seas even during the worst storms were leveled during Tropical Storm Fay in 2008, Hurricane Irene in 2011 and Hurricane Sandy in 2012. A stretch of beachfront railroad track built by NASA in the early 1960s that runs parallel to the shoreline has been topped by waves repeatedly during recent storms. …The research came about after NASA partnered with the U.S. Geological Survey and UF to figure out why chronic erosion was happening along a roughly 6-mile stretch of beach between launch pads 39A and 39B – the ones used for Space Shuttle and Apollo missions. The problem had been occurring for years but seemed to be growing worse, beginning with the spate of hurricanes that struck Florida in 2004. Jaeger said he, Adams and doctoral students Shaun Kline and Rich Mackenzie determined the cause was a gap in a near-shore sandbar that funnels the sea toward that section of beach. … |
| 535 | …Hopefully this research will aid the U.S. Army Corps of Engineers in identifying levees that need repair and assessing the lifespan of structures like the MRGO that are not planned for upkeep. Using fieldwork conducted in the Florida Panhandle, Straub and his fellow researchers were able to better understand the process of seepage erosion, which occurs when the re-emergence of groundwater at the surface shapes the Earth’s topography. The paper suggests that the velocity at which channel heads advance is proportional to the flux of groundwater to the heads. The researchers used field observations and numerical modeling to come up with the theory. … |
Retrieved Docs: [19, 230]
Doc 19: …Based on the ACE projection, combined with the above-average numbers of named storms and hurricanes, the likelihood of an above-normal Atlantic hurricane season has increased to 85%, with only a 10% chance of a near-normal season and a 5% chance of a below-normal season …, said Gerry Bell, Ph.D., lead seasonal hurricane forecaster at NOAA’s Climate Prediction Center …
Retrieved Docs: [230, 273]
Doc 230: …NOAA is predicting a 70 percent likelihood of 6 to 11 named storms, of which 3 to 6 could become hurricanes …While a below-normal season is likely (70 percent), there is also a 20 percent chance of a near-normal season, and a 10 percent chance of an above-normal season …
Retrieved Docs: [230, 636]
Doc 230: …NOAA is predicting a 70 percent likelihood of 6 to 11 named storms, of which 3 to 6 could become hurricanes …While a below-normal season is likely (70 percent), there is also a 20 percent chance of a near-normal season, and a 10 percent chance of an above-normal season …
Retrieved Docs: [583, 275]
Doc 583: …The Climate Prediction Center outlook calls for considerable activity with a 65 percent probability of an above normal season and a 25 percent probability of a near normal season. This means there is a 90 percent chance of a near or above normal season. …
5.4. Ablation Study
Model Components. As shown in Section 3.3, Hypercube-RAG combines two retrieval strategies: sparse exact matching and dense embedding retrieval. Section 3.4 introduces the document ranking algorithm used to filter and prioritize the retrieved results. To study their effectiveness, we select the SciDCC dataset and conduct an ablation study by removing each of them from Hypercube-RAG, which are represented as No-Sparse, No-Dense, and No-Ranking, respectively. Figure 6 shows that the full version of Hypercube-RAG consistently outperforms all other variants, verifying positive contributions of each constituent. More specifically, it justifies the benefit of coupling these two strategies for scientific question-answering tasks, and the necessity of the ranking algorithm to filter possible “noisy” documents.
Hypercube Dimensions. We also conduct an ablation study on the hypercube dimensions: Location, Event, Date, Organization, Person, and Theme. To study their effectiveness, we remove each of the dimensions, which are represented as No-Location, No-Event, No-Date, No-Organization, No-Person, and No-Theme, respectively. Table 8 shows that the full version of Hypercube-RAG consistently outperforms all other variants, verifying positive contributions of each dimension. We observe that the performance drops the most when removing the Location and Theme dimensions, justifying their high importance for theme-specific QA tasks in geoscience. We claim that these dimensions in a hypercube could be changed over datasets with different focuses of interest. The hypercube dimensions used for other datasets are provided in Appendix A.
| Dimension | Hurricane | Geography | Aging Dam |
| No-Location | 78.7 | 65.9 | 77.5 |
| No-Event | 79.6 | 75.9 | 80.0 |
| No-Date | 79.7 | 74.5 | 80.4 |
| No-Organization | 79.6 | 74.6 | 80.2 |
| No-Person | 80.2 | 73.2 | 81.4 |
| No-Theme | 77.9 | 71.7 | 81.1 |
| FULL | 82.5 | 76.8 | 82.6 |
5.5. Parameter Study
Hypercube-RAG combines sparse exact match and semantic retrieval strategies, as introduced in Section 3.3. We investigate the effect of the semantic retrieval strategy on the quality of LLM responses under varying similarity thresholds, denoted by . Figure 7 shows that performance improves with increasing threshold values when , suggesting that higher similarity scores are more effective for retrieving semantically relevant documents. However, when , the behavior of semantic RAG begins to gradually degrade to the sparse exact matching strategy, resulting in a decline in performance. This is primarily due to overly strict matching criteria, which reduce the number of retrieved documents and thus limit the available context.
5.6. Case Study
An additional case study in Figure 5 presents that Sparse BM25, Semantic-RAG (e5), and HippoRAG 2 fail to generate the correct answer due to inaccurate document retrieval. These methods tend to prioritize returning documents that are semantically similar in the embedding space. For example, semantic embedding methods tend to retrieve those documents that largely mention above-normal, near-normal, and below-normal hurricane season. However, these common terminologies may overshadow more specific, less frequent, but critical information, such as the location “Atlantic” and the organization “Climate Prediction Center”. In contrast, our Hypercube-RAG returns the correct answer. It accurately identifies and retrieves documents containing key elements of the query, including “Atlantic”, “hurricane season”, and “Climate Prediction Center”, corresponding to the hypercube dimensions, LOCATON, THEME, and ORGANIZATION, respectively. This retrieval approach intuitively prioritizes specific contextual information over abstract, ambiguous references such as above-, near-, and below-normal alone.
6. Discussion and Conclusions
In this work, we start by analyzing the strengths and limitations of existing RAG methods for scientific question-answering, which tend to perform poorly in either accuracy, efficiency, or explainability of the retrieval process. To retain their advantages while addressing key challenges, we introduce Hypercube, a novel structure representing knowledge in a compact, multi-dimensional space. Cube cell(s) are associated with compact, fine-grained labels extracted from the documents using LLMs.
Built on the hypercube, Hypercube-RAG is proposed to innovate the knowledge retrieval based on those cube labels. The cube label-based retrieval integrates sparse lexical and dense semantic strategies, making it accurate, efficient, and inherently explainable. We conduct experiments on three datasets across different domains comprehensively, and report the results using four retrieval metrics and two QA metrics, demonstrating that our method consistently outperforms existing baselines in both retrieval and question-answering accuracy. Additionally, we reveal a higher retrieval efficiency of Hypercube-RAG than existing approaches, and verify the explainable retrieval process. Overall, we conclude that Hypercube-RAG achieves high accuracy and efficiency, while providing interpretable provenance for retrieval.
Acknowledgements.
This work is supported by the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE), which is funded by the National Science Foundation (NSF) under award number: 2118329. Any opinions, findings, conclusions, or recommendations expressed herein are those of the authors and do not necessarily represent the views of NSF.References
- (1)
- Asai et al. (2023) Akari Asai, Zeqiu Wu, Yizhong Wang, Avirup Sil, and Hannaneh Hajishirzi. 2023. Self-rag: Learning to retrieve, generate, and critique through self-reflection. In The Twelfth International Conference on Learning Representations.
- Augenstein et al. (2024) Isabelle Augenstein, Timothy Baldwin, Meeyoung Cha, Tanmoy Chakraborty, Giovanni Luca Ciampaglia, David Corney, Renee DiResta, Emilio Ferrara, Scott Hale, Alon Halevy, et al. 2024. Factuality challenges in the era of large language models and opportunities for fact-checking. Nature Machine Intelligence 6, 8 (2024), 852–863.
- Chandrasekaran and Mago (2021) Dhivya Chandrasekaran and Vijay Mago. 2021. Evolution of semantic similarity—a survey. Acm Computing Surveys (Csur) 54, 2 (2021), 1–37.
- Cui et al. (2025) Hejie Cui, Alyssa Unell, Bowen Chen, Jason Alan Fries, Emily Alsentzer, Sanmi Koyejo, and Nigam Shah. 2025. TIMER: Temporal Instruction Modeling and Evaluation for Longitudinal Clinical Records. In Will Synthetic Data Finally Solve the Data Access Problem? https://openreview.net/forum?id=uBCAtA6M73
- Devezas and Nunes (2021) José Devezas and Sérgio Nunes. 2021. A review of graph-based models for entity-oriented search. SN Computer Science 2, 6 (2021), 437.
- Edge et al. (2024) Darren Edge, Ha Trinh, Newman Cheng, Joshua Bradley, Alex Chao, Apurva Mody, Steven Truitt, Dasha Metropolitansky, Robert Osazuwa Ness, and Jonathan Larson. 2024. From local to global: A graph rag approach to query-focused summarization. arXiv preprint arXiv:2404.16130 (2024).
- Feng et al. (2015) Wei Feng, Chao Zhang, Wei Zhang, Jiawei Han, Jianyong Wang, Charu Aggarwal, and Jianbin Huang. 2015. STREAMCUBE: Hierarchical spatio-temporal hashtag clustering for event exploration over the Twitter stream. In 2015 IEEE 31st international conference on data engineering. IEEE, 1561–1572.
- Guo et al. (2024) Zirui Guo, Lianghao Xia, Yanhua Yu, Tu Ao, and Chao Huang. 2024. Lightrag: Simple and fast retrieval-augmented generation. arXiv preprint arXiv:2410.05779 (2024).
- Gutiérrez et al. (2024) Bernal Jiménez Gutiérrez, Yiheng Shu, Yu Gu, Michihiro Yasunaga, and Yu Su. 2024. Hipporag: Neurobiologically inspired long-term memory for large language models. In The Thirty-eighth Annual Conference on Neural Information Processing Systems.
- Gutiérrez et al. (2025) Bernal Jiménez Gutiérrez, Yiheng Shu, Weijian Qi, Sizhe Zhou, and Yu Su. 2025. From RAG to Memory: Non-Parametric Continual Learning for Large Language Models. arXiv preprint arXiv:2502.14802 (2025).
- Han et al. (2024) Haoyu Han, Yu Wang, Harry Shomer, Kai Guo, Jiayuan Ding, Yongjia Lei, Mahantesh Halappanavar, Ryan A Rossi, Subhabrata Mukherjee, Xianfeng Tang, et al. 2024. Retrieval-augmented generation with graphs (graphrag). arXiv preprint arXiv:2501.00309 (2024).
- Huang et al. (2025) Lei Huang, Weijiang Yu, Weitao Ma, Weihong Zhong, Zhangyin Feng, Haotian Wang, Qianglong Chen, Weihua Peng, Xiaocheng Feng, Bing Qin, et al. 2025. A survey on hallucination in large language models: Principles, taxonomy, challenges, and open questions. ACM Transactions on Information Systems 43, 2 (2025), 1–55.
- Izacard et al. (2021) Gautier Izacard, Mathilde Caron, Lucas Hosseini, Sebastian Riedel, Piotr Bojanowski, Armand Joulin, and Edouard Grave. 2021. Unsupervised dense information retrieval with contrastive learning. arXiv preprint arXiv:2112.09118 (2021).
- Izacard et al. (2022) Gautier Izacard, Mathilde Caron, Lucas Hosseini, Sebastian Riedel, Piotr Bojanowski, Armand Joulin, and Edouard Grave. 2022. Unsupervised Dense Information Retrieval with Contrastive Learning. Transactions on Machine Learning Research (2022). https://openreview.net/forum?id=jKN1pXi7b0
- Ji et al. (2019) Xiaonan Ji, Han-Wei Shen, Alan Ritter, Raghu Machiraju, and Po-Yin Yen. 2019. Visual exploration of neural document embedding in information retrieval: Semantics and feature selection. IEEE transactions on visualization and computer graphics 25, 6 (2019), 2181–2192.
- Jin et al. (2025) Bowen Jin, Jinsung Yoon, Jiawei Han, and Sercan O Arik. 2025. Long-Context LLMs Meet RAG: Overcoming Challenges for Long Inputs in RAG. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=oU3tpaR8fm
- Kadhim (2019) Ammar Ismael Kadhim. 2019. Term weighting for feature extraction on Twitter: A comparison between BM25 and TF-IDF. In 2019 international conference on advanced science and engineering (ICOASE). IEEE, 124–128.
- Kang et al. (2024) SeongKu Kang, Shivam Agarwal, Bowen Jin, Dongha Lee, Hwanjo Yu, and Jiawei Han. 2024. Improving retrieval in theme-specific applications using a corpus topical taxonomy. In Proceedings of the ACM Web Conference 2024. 1497–1508.
- Karpukhin et al. (2020) Vladimir Karpukhin, Barlas Oguz, Sewon Min, Patrick SH Lewis, Ledell Wu, Sergey Edunov, Danqi Chen, and Wen-tau Yih. 2020. Dense Passage Retrieval for Open-Domain Question Answering.. In EMNLP (1). 6769–6781.
- Krishna et al. (2021) Kalpesh Krishna, Aurko Roy, and Mohit Iyyer. 2021. Hurdles to Progress in Long-form Question Answering. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (Eds.). Association for Computational Linguistics, Online, 4940–4957. doi:10.18653/v1/2021.naacl-main.393
- Lee et al. (2025) Chankyu Lee, Rajarshi Roy, Mengyao Xu, Jonathan Raiman, Mohammad Shoeybi, Bryan Catanzaro, and Wei Ping. 2025. NV-Embed: Improved Techniques for Training LLMs as Generalist Embedding Models. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=lgsyLSsDRe
- Lee et al. (2021) Kenton Lee, Ming-Wei Chang, and Kristina Toutanova. 2021. Learning Dense Representations of Phrases at Scale. In ACL.
- Lin et al. (2023) Xi Victoria Lin, Xilun Chen, Mingda Chen, Weijia Shi, Maria Lomeli, Richard James, Pedro Rodriguez, Jacob Kahn, Gergely Szilvasy, Mike Lewis, et al. 2023. Ra-dit: Retrieval-augmented dual instruction tuning. In The Twelfth International Conference on Learning Representations.
- Liu et al. (2024) Nelson F Liu, Kevin Lin, John Hewitt, Ashwin Paranjape, Michele Bevilacqua, Fabio Petroni, and Percy Liang. 2024. Lost in the middle: How language models use long contexts. Transactions of the Association for Computational Linguistics 12 (2024), 157–173.
- Maillard et al. (2021) Jean Maillard, Barlas Oguz, and Wen-tau Yih. 2021. Multi-Faith Retrieval for Question Answering. In NAACL.
- Mishra and Mittal (2021) Prakamya Mishra and Rohan Mittal. 2021. Neuralnere: Neural named entity relationship extraction for end-to-end climate change knowledge graph construction. In Tackling climate change with machine learning workshop at ICML.
- Peng et al. (2024) Boci Peng, Yun Zhu, Yongchao Liu, Xiaohe Bo, Haizhou Shi, Chuntao Hong, Yan Zhang, and Siliang Tang. 2024. Graph retrieval-augmented generation: A survey. arXiv preprint arXiv:2408.08921 (2024).
- Pipitone and Alami (2024) Nicholas Pipitone and Ghita Houir Alami. 2024. Legalbench-rag: A benchmark for retrieval-augmented generation in the legal domain. arXiv preprint arXiv:2408.10343 (2024).
- Pujari et al. (2023) Tejaskumar Pujari, Anil Kumar Pakina, and Anshul Goel. 2023. Explainable AI and governance: Enhancing transparency and policy frameworks through retrieval-augmented generation (RAG). IOSR Journal of Computer Engineering (2023).
- Robertson and Walker (1994) Stephen E Robertson and Steve Walker. 1994. Some simple effective approximations to the 2-poisson model for probabilistic weighted retrieval. In SIGIR’94: Proceedings of the Seventeenth Annual International ACM-SIGIR Conference on Research and Development in Information Retrieval, organised by Dublin City University. Springer, 232–241.
- Tao et al. (2018) Fangbo Tao, Chao Zhang, Xiusi Chen, Meng Jiang, Tim Hanratty, Lance Kaplan, and Jiawei Han. 2018. Doc2cube: Allocating documents to text cube without labeled data. In 2018 IEEE International Conference on Data Mining (ICDM). IEEE, 1260–1265.
- Trotman et al. (2014) Andrew Trotman, Antti Puurula, and Blake Burgess. 2014. Improvements to BM25 and language models examined. In Proceedings of the 19th Australasian Document Computing Symposium. 58–65.
- Wadden et al. (2020) David Wadden, Shanchuan Lin, Kyle Lo, Lucy Lu Wang, Madeleine van Zuylen, Arman Cohan, and Hannaneh Hajishirzi. 2020. Fact or fiction: Verifying scientific claims. arXiv preprint arXiv:2004.14974 (2020).
- Wang et al. (2025) Gelan Wang, Yu Liu, Shukai Liu, Ling Zhang, and Liqun Yang. 2025. REMFLOW: RAG-enhanced multi-factor rainfall flooding warning in sponge airports via large language model. International Journal of Machine Learning and Cybernetics (2025), 1–21.
- Wang et al. (2024b) Liang Wang, Nan Yang, Xiaolong Huang, Linjun Yang, Rangan Majumder, and Furu Wei. 2024b. Multilingual e5 text embeddings: A technical report. arXiv preprint arXiv:2402.05672 (2024).
- Wang et al. (2024a) Shijie Wang, Jiani Huang, Zhikai Chen, Yu Song, Wenzhuo Tang, Haitao Mao, Wenqi Fan, Hui Liu, Xiaorui Liu, Dawei Yin, et al. 2024a. Graph machine learning in the era of large language models (llms). ACM Transactions on Intelligent Systems and Technology (2024).
- Wang et al. (2022) Yi Wang, Wenhao Yu, Yining Wang, Chenguang Li, Soroush Liu, and Hoifung Poon. 2022. Text Embeddings by Weakly-Supervised Contrastive Pre-training. arXiv preprint arXiv:2212.03533 (2022).
- Wang et al. (2023) Zhaonan Wang, Bowen Jin, Wei Hu, Minhao Jiang, Seungyeon Kang, Zhiyuan Li, Sizhe Zhou, Jiawei Han, and Shaowen Wang. 2023. Geospatial Knowledge Hypercube. In Proceedings of the 31st ACM International Conference on Advances in Geographic Information Systems. 1–4.
- Wellawatte et al. (2025) Geemi P Wellawatte, Huixuan Guo, Magdalena Lederbauer, Anna Borisova, Matthew Hart, Marta Brucka, and Philippe Schwaller. 2025. ChemLit-QA: a human evaluated dataset for chemistry RAG tasks. Machine Learning: Science and Technology 6, 2 (2025), 020601.
- Xiong et al. (2021) Lee Xiong, Chenyan Xiong, Ye Li, Kwok-Fung Tang, Jialin Liu, Paul N. Bennett, Junaid Ahmed, and Arnold Overwijk. 2021. Approximate Nearest Neighbor Negative Contrastive Learning for Dense Text Retrieval. In International Conference on Learning Representations. https://openreview.net/forum?id=zeFrfgyZln
- Zhang et al. (2025a) Qinggang Zhang, Shengyuan Chen, Yuanchen Bei, Zheng Yuan, Huachi Zhou, Zijin Hong, Junnan Dong, Hao Chen, Yi Chang, and Xiao Huang. 2025a. A Survey of Graph Retrieval-Augmented Generation for Customized Large Language Models. arXiv preprint arXiv:2501.13958 (2025).
- Zhang et al. (2025b) Yunyi Zhang, Ruozhen Yang, Xueqiang Xu, Rui Li, Jinfeng Xiao, Jiaming Shen, and Jiawei Han. 2025b. Teleclass: Taxonomy enrichment and llm-enhanced hierarchical text classification with minimal supervision. In Proceedings of the ACM on Web Conference 2025. 2032–2042.
Appendix
Appendix A Hypercube
A.1. Hypercube Design and Construction
Comprehensively understanding the corpus is critical for pre-defining hypercube dimensions, because these dimensions structurally represent the knowledge (i.e., documents) in the multi-dimensional space. We follow these steps to determine the hypercube dimensions from the corpus.
A.2. Hypercube Structure (Dimensions)
Table 9 presents the hypercube dimensions used for different datasets.
| Datasets | Hypercube Dimensions |
| SciFact | “date”, ‘location”, “person”, “quantity”, “organizations_research_initiatives”, “medicine_health”, “genetics_biology”, “immunology_neuroscience”, “pharmacology” |
| LegalBench | “date”, “organization”, “quantity”, “location”, “person”, “company”, “money_finance”, “relationship”, “law_agreement_regulation”, “agreement_information” |
| SciDCC (Hurricane) | “location”, “person”, “event”, “organization”, “theme”, “date” |
| SciDCC (Geography) | “location”, “person”, “event”, “organization”, “theme”, “date” |
| SciDCC (Aging_Dam) | “dam”, “date”, “facility”, “location”, “organization”, “person”, “event”, “quant”, ’watershed’ |
Appendix B Prompt Templates
1. Comprehend the given question: understand what the question asks, how to answer it step by step, and all concepts, aspects, or directions that are relevant to each step.
2. Compose queries to retrieve documents for answering the question: each document is indexed by the entities or phrases occurred inside and those entities or phrases lie within the following dimensions: {dimensions}.
For each of the above dimensions, synthesize queries that are informative, self-complete, and mostly likely to retrieve target documents for answering the question. Note that each of your queries should be an entity or a short phrase and its associated dimension.
Example Input: “How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay?”
Example Output:
1. query_dimension: ‘location’; query_content: ‘Atlantic’;
2. query_dimension: ‘event’; query_content: ‘Tropical Storm Fay’;
3. query_dimension: ’theme’; query_content: ‘rainfall’
Question: {question},
Retrieved documents: {retrieved document},
Output requirements:
- If the query asks the quantitative analysis, such as starting with ”How many”, ”How much”, ”How much greater”, ”How wide”, ”What percentage”, ”How far”, ”How long”, ”How old”, ”What portion”, ”What depth”, please directly output the quantitative answer as short as possible.
- If the query asks specific information, such as ”what percentage”, ”what likelihood”, ”which years”, ”who”, please directly output the answer without explanation or other information.
Question: {question},
Gold Answer: {gold_answer},
Predicted Answer: {predicted_answer}.
Evaluate criteria:
1. Correctness (0 or 1):
- Score 1 if the predicted answer is generally accurate and aligns with the key points in the reference answer.
- Score 0 if it has factual errors or misrepresents key information.
2. Completeness (0 or 1):
- Score 1 if the predicted answer covers the main points present in the reference answer.
- Score 0 if it misses essential information or fails to address the core of the question.
Output your evaluation in the following JSON format:
{“correctness”: int, “completeness”: int, “explanation”: your brief explanation}
Appendix C Discussion
In the following, we discuss different scenarios of hypercube usage: simple query, long query with multiple themes, and diverse query with shifted themes.
C.1. Case 1: simple query
How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay? The relevant documents can be retrieved from one cube cell in a hypercube ¡Melbourne Beach, rainfall, Tropical Storm Fay ¿.
C.2. Case 2: long query with multiple themes
In cases where a query is very diverse, they may need to access multiple cube cells such that the query information can be covered as much as possible. For example, the diverse query could be “What consequences were caused by Tropical Storm Fay, such as how much rainfall Melbourne Beach and Palm Bay in Florida received, and how much monetary losses were caused?”. Those cube cells needed could be ¡Melbourne Beach, rainfall, Tropical Storm Fay ¿, ¡Palm Beach, rainfall, Tropical Storm Fay ¿, ¡Melbourne Beach, monetary losses, Tropical Storm Fay ¿, ¡Palm Beach, monetary losses, Tropical Storm Fay¿.
C.3. Case 3: diverse query with shifted themes
Given a more complicated query including shifted topics, e.g., “How did the monetary loss caused by Tropical Storm Fay impact the industrial layoff in Miami Downtown?”, cube cells across two hypercubes are needed ¡Miami Downtown, Industrial Layoff, Tropical Storm Fay ¿ and ¡Miami Downtown, Monetary loss, Tropical Storm Fay ¿ to get more relevant documents. These cross-theme queries are usually complex and need more dedicated hypercube design, which is considered for future exploration.